OpenStudy (anonymous):

simplify: cot(theta)cos^2(theta)-tan(theta)sin^2(theta)+tan(theta)

OpenStudy (anonymous):

\[ \cot(\theta) \cos^2 (\theta)-\tan (\theta) \sin^2(\theta)+\tan(\theta) \]

OpenStudy (anonymous):

Bream them into trig ratios and see what you get.

OpenStudy (anonymous):

trig ratios?

OpenStudy (anonymous):

tan is sin over cos cot is cos over sin ...

OpenStudy (anonymous):

that one's tough.

OpenStudy (anonymous):

which chapter in your book is this problem coming out of? I feel like we need to use more complicated trig identities with this one.

OpenStudy (anonymous):

it not from a chapter, it's from a worksheet

OpenStudy (anonymous):

do you have answers?

OpenStudy (anonymous):

nope